The prior probabilities for events A1 and A2 are P(A1) = .50 and P(A2) = .50. It is also known that P(A1 A2) = 0. Suppose P(B | A1) = .10 and P(B | A2) = .04. Are events A1 and A2 mutually exclusive? Compute P(A1 B) (to 4 decimals). Compute P(A2 B) (to 4 decimals). Compute P(B) (to 4 decimals). Apply Bayes' theorem to compute P(A1 | B) (to 4 decimals). Also apply Bayes' theorem to compute P(A2 | B) (to 4 decimals).
Answer #1

Ans:

P(A1 B)=P(B | A1)*P(A1)=0.10*0.50=0.0500

P(A2 B)= P(B | A2)*P(A2)=0.04*0.50=0.0200

P(B)=P(A1 B)+P(A2 B)=0.0500+0.0200=0.0700

Apply Bayes' theorem:

P(A1 | B)=P(B | A1)*P(A1)/P(B)=0.0500/0.0700=0.7143

P(A2 | B)= P(B | A2)*P(A2)/P(B)=0.0200/0.0700=0.2857